How many anagrams, with or without meaning, does REFEREE have such that: there is no constraint two Rs are separated it contains subword EE . it begins with letter R
Answer #1

1) there are 2 R ; 4 E and 1 F

number of ways to arrange =7!/(2!*4!*1!)=105

2)

number of ways 2 R's are together =6!/(1!*4!*1!) =30

hence number of ways two R's separated=105-30=75

3)

number of ways 4' E's are separated =3!/(2!*1!)=3

hence contains subword EE =105-3=102

4)

number of ways it begins with letter R =6!/(1!*4!*1!) =30